OpenStudy (studytheworld):

10th grade function help? If f(x)=3(x+5)+4/x what is f(a+2). I know you have to plug in (a+2) for each value where x was, but I'm not getting any of my answer choices as solutions? Help!

OpenStudy (studytheworld):

If \[f(x)=3(x+5)+\frac{ 4 }{ x }\] what is \[f(a+2)\] Answer choices: http://prntscr.com/cksdvf

OpenStudy (studytheworld):

@Visarika23

jhonyy9 (jhonyy9):

just you need substitute the (a+2) in place of x so what will get ?

OpenStudy (visarika23):

Is it not B?

OpenStudy (studytheworld):

I don't know, when I calculate I get this long thing that isn't anything close to my choices

jhonyy9 (jhonyy9):

Study i like check your work step by step and i like help you please collaborate - ok. ?

OpenStudy (visarika23):

Ya show ur work

jhonyy9 (jhonyy9):

Study you dont like getting help ? so directly right answers we not can giving bc. this is opposite to the OS rules

OpenStudy (studytheworld):

I'm trying to type it

OpenStudy (studytheworld):

But nvm I can't get it, I don't know how I got my answer. It doesn't really matter since I know it's wrong. I just need help getting the RIGHT answer

OpenStudy (studytheworld):

So I'm not really sure where to start... @Visarika23 @jhonyy9

OpenStudy (visarika23):

See when u put x=a+2 u get 3(a+2+5) + 4/(a+2) 3(a+7) + 4/(a+2)

OpenStudy (retireed):